OpenStudy (anonymous):

If f(x) = 12x – 7, what is f–1(x) ?

OpenStudy (apoorvk):

so y=12x-7 now write 'x' in terms of 'y', and then interchange 'x' and 'y'. the resulting function is the inverse function of f(x).

OpenStudy (anonymous):

i just need the answer

Parth (parthkohli):

f-1(x) is the inverse of f(x) x = 12y - 7 x + 7 = 12y \(x + 7 \over 12\) = y y is the inverse.

Parth (parthkohli):

Therefore, \(\Large \color{purple}{\rightarrow f^{-1}(x) = {x + 7 \over 12} }\)

OpenStudy (anonymous):

so x-12 over 7?

OpenStudy (anonymous):

or is x+7/12 the answer

OpenStudy (apoorvk):

Sorry, but we don't just give the answers here. This is a tutoring website, not an answering one. It'd rather help you if you could learn the method - rather than just 'copying' the answers.

OpenStudy (anonymous):

i have one question left on my paper and im frustrated and just want to get it over with